A 0.22 kg piece of silly putty is thrown at a 1.2 kg wooden panel that is initially at the rest. The velocity of the silly putty before the collision was 18.5 m/s. The silly putty sticks to the wooden panel and they move as one after the collision. What is the velocity of the combined objects after the collision.


1
Expert's answer
2020-11-05T10:33:39-0500

"m_1v_1=(m_1+m_2)v_2\to v_2=\frac{m_1v_1}{m_1+m_2}=\frac{0.22\cdot 18.5}{0.22+1.2}=2.87(m/s)"